# Crowd-PrefRL: Preference-Based Reward Learning from Crowds

### v1

@@ 3.2 Estimating Preference Labels from a Crowd
For each query, Crowd-PrefRL distills the $M$ preference labels from each crowd member to a single preference label. We discuss two methods for accomplishing this distillation. The first, majority voting, is a comparatively naive approach, while the second, the Spectral Meta Learner (SML) Parisi et al. (2014) is the method we recommend for use with Crowd-PrefRL. Note that our Crowd-PrefRL framework generalizes the crowdsourcing approach in Brown et al. (2019), in which majority voting is used to merge preference labels.
@@
Majority Vote (Baseline). Given the current problem formulation, where we do not know any additional information about the crowd, perhaps the simplest estimate $\hat{\bm{y}}$ is the majority vote (MAJ), which labels each segment pair $x_{k}\in\mathcal{S}$ with the majority preference from the crowd, denoted as $\hat{\bm{y}}_{MAJ}$.
@@
22:         $\mathcal{D}\leftarrow\mathcal{D}\cup\{(A,B,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\hat{\bm{y}}})\}$
@@
24:              Sample batch $\{(A,B,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\hat{\bm{y}}})_{j}\}_{j=1}^{|\mathcal{D}|}\sim\mathcal{D}$
@@
Spectral Meta-Learner. However, it is indeed possible to estimate a crowd label that often outperforms the majority vote in the fully unsupervised case studied here. Under the assumption that users in a crowd make independent errors, Parisi et al. (2014) derived the Spectral Meta-Learner (SML):
$\hat{\bm{y}}_{k}=\operatorname{sign}\left(\sum_{i=1}^{M}f_{i}(x_{k})*\hat{\bm{v}}_{i}\right),$ (2)
where $\hat{\bm{v}}_{i}$ is the lead eigenvector of the covariance matrix obtained from the predictions of a set of binary classifiers on a given query; in our case, these correspond to each user’s binary preference on a given pairwise comparison query. As shown in Parisi et al. (2014), crowd labels derived via the SML are typically more accurate than labels from any individual crowd member or from a majority vote, as unlike other methods such as majority vote, the SML approximates a maximum likelihood estimate of the labels. In addition, Parisi et al. (2014) showed that $\hat{\bm{v}}_{i}$ can rank the users by their estimated error rates, which approximate the ground truth error rates. We denote labels estimated by SML as $\hat{\bm{y}}_{SML}$.
Parisi et al. (2014) also analyze the performance of SML in the presence of a cartel, a small group of users who attempt to steer the overall ensemble solution away from the (unknown) ground truth, and proved that (1) SML is robust to the presence of cartels, and that (2) the weights of the eigenvector $\hat{\bm{v}}_{i}$ can be used to identify the presence of cartels in the crowd. Here, we re-define the notion of a cartel as a minority group who provides feedback that consistently differs from the majority, for example users who provide feedback towards goals that are different than the goals of the majority.
3.3 Crowdsourced Preference Learning for RL
Crowd-PrefRL leverages the crowd-aggregated preference labels $\hat{\bm{y}}=\hat{\bm{y}}_{SML}$ for all instances in $\mathcal{S}$. We denote via $\mathcal{D}$ the dataset of all pairs $\{(x,y)\}$, where $x\in\mathcal{S}$ and $y$ is the corresponding crowd-aggregated label in $\hat{\bm{y}}$. To learn useful behaviors given these aggregated labels, we assume that the majority of users (though not necessarily all users) have a shared understanding of desired behavior, which can be quantified in terms of an underlying reward function. Given the crowd-aggregated labels, our objective is to learn the reward function, and ultimately a policy, that preserves the underlying user preferences Christiano et al. (2017). This is done in two steps: (1) a reward function learning step, where we optimize the learned reward via supervised learning from the preference labels $\hat{\bm{y}}$ over segment pairs generated by the policy $\pi$, and (2) a policy learning step, where the agent performs rollouts in the environment and optimizes its policy against the current reward function estimate. Segment pairs for preference queries are selected according to an entropy-based measure from an ensemble of reward predictors, similarly to Lee et al. (2021). These steps are then repeated for a set number of environment steps. Algorithm 1 shows the algorithmic steps needed for preference-based RL with learned rewards from crowds, and is analogous to Lee et al. (2021) with red text denoting the modifications needed to crowdsource the preference labels rather than obtaining them from a single human.
Similarly to Lee et al. (2021); Christiano et al. (2017), we learn a reward model $\widehat{r}_{\psi}$, modeled as a neural network with parameters $\psi$, that predicts the user preferences as follows:
$\mathcal{L}^{\tt Reward}=-\operatorname*{\mathbb{E}}_{(A,B,y)\sim\mathcal{D}}\Big{[}$ $\mathbbm{1}_{[y=-1]}\log P_{\psi}[B\succ A]+$
$\mathbbm{1}_{[y=1]}\log P_{\psi}[A\succ B]\Big{]},$ (4)
where $\mathbbm{1}_{[\cdot]}$ denotes an indicator variable. Given the learned reward function $\widehat{r}_{\psi}$, a policy $\pi$ can be learned using any RL algorithm. In this work, we use PrefPPO Lee et al. (2021), instantiating Crowd-PrefRL as Crowd-PrefPPO. PrefPPO uses Proximal Policy Optimization (PPO), a state-of-the-art on-policy RL algorithm Schulman et al. (2017). As recommended in Lee et al. (2021), we use on-policy methods to reduce the effects of non-stationarity induced by learning a reward function and policy simultaneously.
4 Experiments
4.1 Experiment Setup
We conduct experiments to evaluate the proposed approach to crowdsourcing human feedback for preference learning. Our experiments conduct RL from human preferences using the PrefPPO implementation in Lee et al. (2021). These experiments use the Walker-walk, Quadruped-walk, and Cheetah-run environments from the DMControl suite Tassa et al. (2018), as two of these environments (Walker-walk and Quadruped-walk) were previously used in Lee et al. (2021). For each environment, we report results for Crowd-PrefPPO (SML preference estimation) and the following comparisons: Crowd-PrefPPO-MAJ (in which the majority vote is used instead of the SML estimate) and Crowd-PrefPPO-Oracle (in which the preference is given to the segment with the highest ground-truth environment reward); in addition, we compare Crowd-PrefPPO’s estimated preference labels to those from the best member of each crowd. We perform 10 total runs of each algorithm comparison in each environment for 4M environment steps. Methods are compared via the mean and standard error over 6 runs, where the top 2 and bottom 2 runs are omitted (measured by episode return at 4M environment steps) to reduce the impact of outliers (similar to the InterQuartile Mean, IQM Agarwal et al. (2021)). All remaining hyper-parameters for our method can be found in the Appendix, and notably, are the same as reported in Lee et al. (2021).
4.2 Results
Figure 2: Difference in MAJ and SML preference prediction error rates for different levels of variability (standard deviation) in user error rates. Values are calculated across 100 randomly-sampled crowds of 7, 11 and 15 simulated users for the Walker-walk environment. The horizontal dashed line at $y=0$ indicates where the MAJ and SML error rates are the same; points above $y=0$ indicate that the SML labels have lower error than the MAJ labels. We see a strong positive correlation, with increasing standard deviations (indicating a more diverse crowd) leading to improved SML label accuracy (lower error rate) compared to MAJ labels. In addition, we see a similar trend across all crowd sizes, suggesting that our method is robust across different sizes of crowds. Red dots indicate where SML outperforms the best crowd member, while blue dots indicate where the best crowd member outperforms SML.
We first study the effect of crowd size on the performance of crowdsourced preference learning by simulating 100 different random crowd configurations at each of three different crowd sizes ($M\in\{7,11,15\}$) according to the Stochastic Preference Model in Equation (1) for the Walker-Walk environment; see Methods for how these diverse crowds are simulated. Figure 2 depicts the difference in errors between the MAJ and SML preference estimates versus the diversity in users’ error rates with respect to the ground-truth reward (quantified via the standard deviation of user error across the crowd) for each simulated crowd size. To calculate these errors for each simulated crowd configuration, we first perform one full training run of Crowd-PrefPPO using Oracle preference labels to generate the full set of segment pairs needed for reward labeling. Then, we use these segment pairs and their ground-truth labels to construct the error rates for all users in the crowd and the error rates corresponding to the MAJ and SML labels. We note that results were similar for the other environments tested, and therefore show results for only the Walker-walk environment due to space constraints. There are several results here that are worth discussing. First, we see a strong positive correlation between the error rates of users in the crowd and the performance of SML compared to MAJ. In particular, as the user error increases (indicating a more diverse crowd), the better the SML labels perform compared to the MAJ labels. This indicates that the SML is effectively filtering out inconsistent preference feedback from users who have higher error rates and aggregating crowd decisions more consistently compared to MAJ. Second, we see that SML outperforms MAJ in nearly all cases ($\approx 90\%$ of points above $y=0$), suggesting that SML captures the crowd preferences more accurately than MAJ. Third, we see that the SML labels have lower errors than the best crowd member’s labels in nearly all cases (red dots). As the crowd size increases, we see fewer instances of SML performing worse than the best member of the crowd (blue dots), while simultaneously having more instances where SML outperforms MAJ (dots above $y=0$ dashed line). Finally, across all crowd sizes, the spread of the difference in MAJ errors and SML errors is fairly consistent (most points fall between $[0,6]\%$), suggesting that SML labels can outperform MAJ labels with most crowd configurations. While this improvement may seem small, it occurs during every reward learning iteration; we believe that this leads to compounding benefits during reward learning, and consequently, during agent policy learning.
Having studied how crowd configuration variability affects the respective prediction errors of the SML and MAJ crowd preference estimators, we next conduct experiments with Crowd-PrefPPO using the crowdsourced preference labels; in these experiments, we aim to determine if the improved label error from SML observed in Figure 2 impacts agent performance in a meaningful way. For this analysis, we select a fixed crowd size of $M=11$ users and sample two different crowd configurations: (1) Figure 3a studies a crowd for which SML outperforms MAJ by approx. $3\%$ in prediction error (following the analysis in Figure 2), and (2) Figure 3b studies a crowd for which SML performs similarly to MAJ in prediction error (similarly from the analysis in Figure 2). In addition, in the bottom row of Figure 3, we plot the error rates of the MAJ and SML label predictions relative to the ground-truth labels during each crowd feedback iteration for all three tested environments (200 queries per iteration for Quadruped-walk and 100 otherwise; see the Appendix for more details).
Figure 4: Plot of ground-truth error rate for each user in the crowd vs. the SML weight vector $\bm{v}_{i}$ for the same two SML configurations shown in Figure 3 in the Walker-walk environment. Across both conditions, we see a significant correlation between the weights generated by the SML and individual crowd member performance measured by the users’ (unknown) error rates, suggesting that users can be ranked accurately in the absence of ground-truth error information about the crowd.
In Figure 3a, we see that Crowd-PrefPPO outperforms Crowd-PrefPPO-MAJ in terms of trajectory return for all tested environments, with Crowd-PrefPPO achieving close to Oracle performance for Walker-walk and Quadruped-walk. In addition, the SML label errors are lower than the MAJ label errors at each feedback iteration in most cases for each environment. The lower SML label errors imply that there is a compounding benefit to using SML labels rather than MAJ labels for reward learning; in early feedback iterations, the learned reward function has lower error and thus is closer to the Oracle reward function compared to MAJ. This improves overall agent training, as (1) the agent receives more accurate rewards for its actions, enabling it to learn better behaviors earlier in the training loop, and (2) future crowd feedback iterations receive “better” segment pairs for the crowd to label, yielding more useful information about the reward function. We speculate that this is why Crowd-PrefPPO’s performance compares fairly well with that of Crowd-PrefPPO-Oracle. We see a similar, but weaker, trend in Figure 3b, where SML is expected to yield a similar label accuracy to MAJ for the given crowd configuration. There is generally less spread in episode returns between SML and MAJ for the three tested environments (top row of Figure 3b), and we see that the crowd feedback errors are closer between the methods during all crowd feedback iterations (bottom row of Figure 3b), indicating only minimal benefits of SML labels over MAJ labels. However, we do see that Crowd-PrefPPO can sometimes perform quite well overall, for example in Quadruped-walk.
Furthermore, we analyze the crowd member ranking derived by SML for the same two conditions shown in Figure 3 and the Walker-walk environment (results were similar for all environments, thus we report only one due to space constraints). While Crowd-PrefPPO calculates the eigenvector $\bm{v}_{i}$ during each feedback iteration, we here instead calculate $\bm{v}_{i}$ using all crowdsourced preference labels for simplicity of visualization. Results are shown in Figure 4. We see that across both conditions, the SML-derived member rankings are significantly correlated to the individual members’ error rates (i.e., higher SML weight values indicate lower error rates), suggesting that even when SML performs similarly to MAJ in label error (Figure 4b), SML can still identify the best members in the crowd. This showcases additional utility of SML aggregation over MAJ aggregation even in the situation where SML provides no benefit in label accuracy compared to MAJ.
Figure 5: Plot of ground-truth error rate vs. the SML weight vector $\bm{v}_{i}$ for a crowd configuration where there are three distinct groups of users: users who provide feedback according to (1) the ground-truth reward $r$ (blue), (2) at random (black), and (3) the opposite of the ground-truth reward, $-r$ (red). We see distinct groupings of users by their feedback category and a strong correlation between their (unknown) error rates and the SML weight vector $\bm{v}_{i}$, suggesting that minority groups of users can be accurately identified.
Finally, we test the ability of SML to detect the presence of minority group feedback based purely on the preference labels. We follow the same analysis used in Figure 4, except that we simulate a crowd of 15 users according to three groups: users who provide noisy preference feedback according to (1) the ground-truth reward $r$ (9 users), (2) at random (3 users), and (3) $-r$, the opposite of the ground-truth (3 users). Results are shown in Figure 5. As expected, we see that MAJ error rates are significantly higher compared to SML rates, since MAJ is more sensitive to the presence of noisy labels. In addition, we see distinct clusters of users based on their feedback category, suggesting that the SML weights $\bm{v}_{i}$ can be used to identify the presence of minority groups of users, as well as the identities of the minority group members. These results suggest that Crowd-PrefRL could be used to cluster users via their objectives and to learn reward functions and corresponding behaviors specific to each user cluster. Such functionality could enable a behavior-based characterization of the objective associated with each user cluster.
5 Discussion
This study aimed to determine how to (1) effectively combine preference feedback from crowds of unknown expertise and reliability in the absence of ground-truth information, (2) identify the “best” users in a given crowd in an unsupervised manner, (3) learn a crowdsourced reward function that is more accurate than the reward learned from any individual user, and (4) identify minority groups based purely on their preference feedback. We propose the Crowd-PrefRL framework and instantiate it as Crowd-PrefPPO. Our evaluation of Crowd-PrefPPO indicates that we can use methods derived from unsupervised ensemble learning (SML, Parisi et al. (2014)) to effectively aggregate user preference feedback across diverse crowds in the absence of any ground-truth information about the users. Furthermore, our evaluation suggests that Crowd-PrefPPO can identify which users in a crowd are likely most reliable and that the reward function learned from crowd-aggregated labels is more accurate than the reward derived from any single user’s feedback (even the best user!). Finally, we show the possibility of detecting clusters of minority feedback in an unsupervised manner, potentially enabling a deeper understanding of model behavior (e.g. detecting the presence of multiple user viewpoints, characterizing the associated objectives, and mitigating effects of malicious user data Siththaranjan, Laidlaw, and Hadfield-Menell (2023); Ouyang et al. (2022)). To the best of our knowledge, this is the first demonstration of (1) learning crowd-level preference labels for preference-based RL via unsupervised ensemble learning methods in the online RL setting and (2) identification of minority groups in the crowd based purely on the crowd’s preference feedback.
Previous work has studied crowdsourcing preference labels for reward learning in the offline RL case  Zhang and Kashima (2023), as opposed to the online RL case we study here. In the offline case, a dataset of preference pairs is already provided a priori, and the goal is to learn a reward function from this static dataset. This limits the expressiveness of the learned reward function to the behaviors contained in the dataset, whereas in our work, the preference pairs for crowd labeling are sampled throughout agent training, and by extension the learned reward function continuously improves with better agent performance and better environment exploration. In addition, the method in  Zhang and Kashima (2023) requires training a separate model (e.g. fully connected network) to infer crowd labels and member reliabilities, requiring additional model training and tuning and increased data requirements (50K preference pairs from 2,500 crowdworkers). In contrast, our approach to inferring crowd labels by the SML is learning and optimization-free: we need no hyperparameters to specify, no model to train, and no optimization procedure to learn a crowd label. Thus, we believe it is more suitable in the online RL case where the number of labeled preference pairs is small compared to the offline RL case due to practical constraints (agent is training during human feedback collection). Our work enables crowds to update the reward function throughout agent training using preference feedback on just a handful of trials and furthermore, allows for the detection of minority feedback groups, which has important modeling applications in various domains.
We showed that SML labels outperform MAJ labels in nearly all cases across different crowd sizes (Figure 2) and different tasks (Figure 3). Calculating SML weights is computationally very fast, only requiring a few lines of code in modern programming languages (e.g. Python). Given the increasing popularity of crowdsourcing platforms for data collection in human-interactive RL studies (e.g. Amazon Mechanical Turk), these results indicate that simple code modifications can result in significant performance gains. For example, Brown et al. (2019) used a crowdsourcing platform for their human subjects data analysis and estimated preference labels for segments by simply taking a majority vote over multiple human users. It would be interesting to apply Crowd-PrefRL to aggregate crowd preferences in such existing studies to determine the extent to which Crowd-PrefRL improves over results from a majority vote or offline RL.
We believe that Crowd-PrefRL could be combined with recent advances that make preference-based RL methods more sample-efficient Lee, Smith, and Abbeel (2021); Park et al. (2022). Improvements to sample efficiency are made primarily at the algorithmic or data level, as opposed to modifying the labels used to learn the reward. For example, PEBBLE Lee, Smith, and Abbeel (2021) combines unsupervised pre-training with off-policy re-labeling of past data to achieve improved sample efficiency compared to PrefPPO. SURF Park et al. (2022) uses semi-supervised learning methods that combine data augmentation approaches with confidence-based pseudo-labels of unlabeled experiences to train the preference predictor. Combining Crowd-PrefRL with these orthogonal advancements could potentially lead to further improvements in RL agent training performance.
Several downsides to our approach are important to discuss. First, each user in the crowd must provide a preference label for each preference query. This results in significant human labeling cost if implemented in real-world studies. Our results indicate that Crowd-PrefPPO can work well in fairly small crowds with relatively small numbers of labels (100 queries per user per reward learning iteration). Future work will focus on the possibility of using Crowd-PrefRL when user feedback is sparse (i.e., not every user provides a preference for every segment pair). In addition, this study uses only simulated human feedback and does not include human-subjects data collection. Rather, we simulate human feedback via the Stochastic Preference Model, which allows for a thorough and systematic evaluation of Crowd-PrefRL and an analysis of its properties in a highly controlled setting. Future work will evaluate this method in a human subjects study.
Appendix
Experiment Setup Hyper-parameters
@@ 5 Discussion
Our experiments use the Walker-walk, Quadruped-walk, and Cheetah-run environments from the DMControl suite Tassa et al. (2018), as two of these environments (Walker-walk and Quadruped-walk) were previously used in Lee et al. (2021). All hyper-parameters for our method can be found in Table 1, and notably, are the same as reported in Lee et al. (2021).

### v2

@@ 4.2 Estimating Preference Labels from a Crowd
For each query, Crowd-PrefRL distills the $M$ preference labels from each crowd member to a single preference label. We discuss two methods for accomplishing this distillation. The first, majority voting, is a comparatively naive approach, while the second, the Spectral Meta Learner (SML) (Parisi et al., 2014), is the method we recommend for use with Crowd-PrefRL.
@@
Majority Vote (Baseline). Given the current problem formulation, in which we lack additional information about the crowd, perhaps the simplest estimate $\hat{\bm{y}}$ is the majority vote (MAJ), which labels each segment pair $x_{k}\in\mathcal{S}$ with the majority preference from the crowd, denoted as $\hat{\bm{y}}_{MAJ}$.
Spectral Meta-Learner. However, it is indeed possible to estimate a crowd label that often outperforms the majority vote in the fully unsupervised case studied here. Under the assumption that users in a crowd make independent errors, Parisi et al. (2014) derived the SML:
$\hat{y}_{k}=\operatorname{sign}\left(\sum_{i=1}^{M}f_{i}(x_{k})*\hat{v}_{i}\right),$ (2)
where $\hat{v}_{i}$ are the values of $\hat{\bm{v}}$, the lead eigenvector of the empirical data covariance matrix of the users’ binary preferences across the set of pairwise comparison queries. As shown in Parisi et al. (2014), crowd labels derived via the SML are typically more accurate than labels from any individual crowd member or from a majority vote, as unlike other methods such as majority vote, the SML approximates a maximum likelihood estimate of the labels. In addition, Parisi et al. (2014) showed that $\hat{\bm{v}}$ can rank the users by their estimated error rates, which approximate the ground truth error rates. We denote the labels estimated by SML as $\hat{\bm{y}}_{SML}$. Parisi et al. (2014) also analyze the performance of SML in the presence of a cartel, a small group of users who attempt to steer the overall ensemble solution away from the (unknown) ground truth, and proved that (1) SML is robust to the presence of cartels, and that (2) the weights of the eigenvector $\hat{\bm{v}}$ can be used to identify the presence of cartels in the crowd. Here, we redefine the notion of a cartel as a minority group of users who provide feedback that consistently differs from that of the majority, for example users who provide feedback towards goals that are different from the goals of the majority. Notably, to our knowledge, the SML has not previously been applied toward preference-based learning or RL.
4.3 Crowdsourced Preference Learning for RL
Crowd-PrefRL (see Algorithm 1 in the Supplemental Material) leverages the crowd-aggregated preference labels $\hat{\bm{y}}=\hat{\bm{y}}_{SML}$ for all instances in $\mathcal{S}$. We denote via $\mathcal{D}$ the dataset of all pairs $\{(x,y)\}$, where $x\in\mathcal{S}$ and $y$ is the corresponding crowd-aggregated label in $\hat{\bm{y}}$. To learn useful behaviors given these aggregated labels, we assume that the majority of users (though not necessarily all users) have a shared understanding of desired behavior, which can be quantified in terms of an underlying reward function. This is done by iterating between the following two steps: (1) a reward function learning step, where we optimize the learned reward via supervised learning on the preference labels $\hat{\bm{y}}$ given over behaviors generated by the policy $\pi$, and (2) a policy learning step, where the agent performs environment rollouts and optimizes its policy against the current reward function estimate. Segment pairs for preference queries are selected according to the disagreement among an ensemble of reward predictors, similarly to Lee et al. (2021a). We then learn a reward model $\widehat{r}_{\psi}$, modeled as a neural network with parameters $\psi$, that predicts the user preferences as follows:
$\mathcal{L}^{\tt Reward}=-\operatorname*{\mathbb{E}}_{(A,B,y)\sim\mathcal{D}}\Big{[}$ $\mathbbm{1}_{[y=-1]}\log P_{\psi}[B\succ A]+\mathbbm{1}_{[y=1]}\log P_{\psi}[A\succ B]\Big{]},$ (4)
where $\mathbbm{1}_{[\cdot]}$ denotes an indicator variable. Given the learned reward function $\widehat{r}_{\psi}$, a policy $\pi$ can be learned using any RL algorithm. In this work, we use PrefPPO (Lee et al., 2021a), instantiating Crowd-PrefRL as Crowd-PrefPPO. PrefPPO uses Proximal Policy Optimization (PPO), a state-of-the-art on-policy RL algorithm (Schulman et al., 2017). As recommended in Lee et al. (2021a), we use on-policy methods to reduce the effects of non-stationarity induced by learning a reward function and policy simultaneously.
5 Experiments
Figure 2: Difference in MAJ and SML preference prediction error rates for different levels of variability (standard deviation) in user error rates. Values are calculated across 100 randomly-sampled crowds of 7, 11 and 15 simulated users for the Walker-walk environment. The horizontal dashed line at $y=0$ indicates where the MAJ and SML error rates are the same; points above $y=0$ indicate that the SML labels have lower error than the MAJ labels. Red dots indicate where SML outperforms the best crowd member; blue indicates where the best crowd member outperforms SML.
5.1 Experiment Setup
We conduct experiments to evaluate the proposed approach to crowdsourcing human feedback for preference learning. Our experiments conduct RL from human preferences using the PrefPPO implementation in Lee et al. (2021a). These experiments use the Walker-walk, Quadruped-walk, and Cheetah-run environments from the DMControl suite (Tassa et al., 2018). For each environment, we report results for Crowd-PrefPPO-SML (SML preference estimation) and the following comparisons: Crowd-PrefPPO-MAJ (in which the majority vote is used instead of the SML estimate) and Crowd-PrefPPO-Oracle (in which the preference is given to the segment with the highest ground-truth environment reward); in addition, we compare Crowd-PrefPPO’s estimated preference labels to those from the best member of each crowd. We perform 10 total runs of each algorithm comparison in each environment for 4M environment steps. Methods are compared via the mean and standard error over 6 runs, where the top 2 and bottom 2 runs are omitted (measured by episode return at 4M environment steps) to reduce the impact of outliers (similar to the InterQuartile Mean, IQM (Agarwal et al., 2021)). All remaining hyperparameters are the same as reported in Lee et al. (2021a) and listed in the Supplemental Materials.
5.2 Results
We first study the effect of crowd size on the performance of crowdsourced preference learning by simulating 100 different random crowd configurations at each of three different crowd sizes ($M\in\{7,11,15\}$) according to the Stochastic Preference Model in Equation (1) for the Walker-Walk environment. Figure 2 depicts the difference in errors between the MAJ and SML preference estimates versus the diversity in users’ error rates with respect to the ground-truth reward (quantified via the standard deviation of user error across the crowd) for each simulated crowd size. To calculate these errors for each simulated crowd, we ran Crowd-PrefPPO using Oracle preference labels to generate sets of segment pairs for preference labeling and reward learning. Then, we use these segment pairs and their ground-truth labels to construct the error rates for all individual users in the crowd and the error rates corresponding to the MAJ and SML aggregated labels. Since results were similar for the other environments tested, we show results for only the Walker-walk environment due to space constraints. We see a positive correlation between the error rates of users in the crowd and the performance of SML compared to MAJ. In particular, as the user variability increases, the better the SML labels perform compared to the MAJ labels, indicating that SML is filtering out inconsistent preference feedback from users with higher error rates, and SML is aggregating crowd decisions more consistently than MAJ. We also see that SML outperforms MAJ in nearly all cases ($\approx 90\%$ of points above $y=0$), suggesting that SML captures the crowd preferences more accurately than MAJ. In addition, the SML labels have lower errors than the best crowd member’s labels in nearly all cases (red dots). Finally, across all crowd sizes, the spread of the difference in MAJ errors and SML errors is fairly consistent (most points fall between $[0,4]\%$), suggesting that SML labels can outperform MAJ labels with most crowd configurations. While this improvement may seem small, it occurs during every reward learning iteration; we believe that this leads to compounding benefits during reward learning, and consequently, during agent policy learning.
We next conduct experiments with Crowd-PrefPPO to determine if the improved label error from SML observed in Figure 2 impacts agent performance in a meaningful way. For this analysis, we select a fixed crowd size of $M=15$ users and sample two different crowd configurations: (1) Figure 3a studies a crowd for which SML outperforms MAJ by $\approx 4\%$ in prediction error (following the analysis in Figure 2), and (2) Figure 3b studies a crowd for which SML performs similarly to MAJ in prediction error (similarly from the analysis in Figure 2). In addition, in the bottom row of Figure 3, we plot the error rates of the MAJ and SML label predictions relative to the ground-truth labels during each crowd feedback iteration for all three tested environments. We see that Crowd-PrefPPO with SML label aggregation (Crowd-PrefPPO-SML) outperforms Crowd-PrefPPO with MAJ labels (Crowd-PrefPPO-MAJ) in terms of trajectory return for all tested environments, with Crowd-PrefPPO achieving close to Oracle performance for Walker-walk and Quadruped-walk. In addition, the SML label errors are lower than the MAJ label errors at each feedback iteration in most cases for each environment. The lower SML label errors imply that there is a compounding benefit to using SML labels rather than MAJ labels for reward learning; in early feedback iterations, the learned reward has lower error and thus is closer to the Oracle reward function compared to MAJ, improving agent training. We see a similar, but weaker, trend in Figure 3b, where SML is expected to yield a similar label accuracy to MAJ for the given crowd configuration. There is generally less spread in episode returns between SML and MAJ for the three tested environments (Figure 3b, top), and we see that the crowd feedback errors are closer between the methods during all crowd feedback iterations (Figure 3b, bottom), indicating only minimal benefits of SML labels over MAJ labels.
Figure 4: SML weights and agent performance in MO-Hopper. (a) 25K training steps. Left: Scatter plot of crowd SML weights $\hat{v}_{i}$ with corresponding user error rates. Right: scatter plot of returns of 100 rollout trajectories with two reward functions, (1) Forward and (2) Vertical Height. (b) Similar to (a) with 100K training steps. At 25K training steps (a), we observe a reliable separation between the minority and majority groups and weakly negatively correlated returns in the two reward functions, indicating that the agent has not yet learned to optimize both objectives. In contrast, at 100K training steps (b), the agent has learned to satisfy both sets of users, as indicated by the large positive correlation in returns, and thus the SML weights no longer separate into two groups.
Finally, we conduct an experiment to test the ability of SML to detect the presence of minority group feedback based purely on the crowd preference labels. For this experiment, we use the MO-Hopper environment found in Multi-Objective Gymnasium (Felten et al., 2023). This environment encodes two reward functions according to two different objectives: (1) forward velocity, which encourages the agent to run forward and (2) vertical height, which encourages the agent to jump high. We simulate a crowd of 15 users, 11 of whom assign preference feedback according to Objective 1 (forward velocity) while the remainder assign feedback according to Objective 2 (vertical height). We aim to determine if the SML weights can identify the presence of minority groups within the crowd. To obtain the initial set of trajectories for crowd labeling, we first train an agent via PPO using stable-baselines3 (Raffin et al., 2021) on a linearized reward that averages the two reward functions, and evaluate the learned model after both 25K and 100K training frames. At both of these checkpoints, we collect 100 agent rollout trajectories, sample 100 trajectory pairs for crowd labeling, and calculate the SML labels and each user’s SML weight $\hat{v}_{i}$. Results are shown in Figure 4. First, in Figure 4a (left), we see that at 25K training steps, there is a separation of the minority (orange) and majority (blue) crowd SML weights, with an approximate threshold at $\hat{v}_{i}=0$ separating the two groups. We also see (Figure 4a, right) that the trajectory returns are weakly negatively correlated across the two objectives, indicating that the agent has not yet learned to optimize both objectives simultaneously (e.g. preferring to run forward instead of jump for some trajectories, and vice-versa). At 100K training steps (Figure 4b), we see a different result: the agent now can optimize behaviors according to both objectives (as indicated by high positive correlation in the two returns), while the SML weights can no longer separate the presence of minority and majority groups. This is because the agent behaviors now satisfy the objectives of both the minority and majority groups, so that even though the two groups assign feedback according to different reward functions, there is no longer any disagreement in crowd preferences.
Figure 5: Unsupervised clustering of users based purely on crowd feedback. We consider a crowd of 150 workers with a 110/40 majority/minority split in the MO-Hopper environment. (a) For 25K agent training steps, we see (left) a histogram of SML weights $\hat{v}_{i}$ overlaid with a fitted GMM, (middle) the BIC values used to automatically infer the number of clusters, and (right) a scatter plot of SML weights vs. ground-truth error showing the accuracy of the inferred grouping. Colors indicate the ground truth grouping (blue for majority, orange for minority). (b) Similar to (a) with 100K training steps. We see that for the agent at 25K training steps (a), a GMM with two Gaussian components reliably models the two groups using only the SML weights $\hat{v}_{i}$, automatically inferring the number of clusters and performing unsupervised cluster assignment. At 100K training frames (b), the agent can satisfy both sets of users, and so the cluster analysis no longer infers any minority.
[…]
6 Discussion
This study proposes the Crowd-PrefRL framework for training AI agents to integrate feedback from many users. Our evaluation of Crowd-PrefPPO indicates that we can use methods derived from unsupervised ensemble learning (SML, Parisi et al. (2014)) to effectively aggregate user preference feedback across diverse crowds to learn RL policies without any ground-truth user information. Furthermore, our evaluation suggests that Crowd-PrefPPO can learn a reward from crowd-aggregated labels that is more accurate than that derived from any single user’s feedback. Finally, we show the possibility of detecting minority feedback clusters in an unsupervised manner, potentially enabling a deeper understanding of model behavior, e.g. via detecting multiple user viewpoints, characterizing the associated objectives, and mitigating the effects of malicious user data.
We showed that SML labels outperform MAJ labels in nearly all cases across different crowd sizes (Figure 2) and different tasks (Figure 3). Calculating SML weights is computationally very fast, only requiring a few lines of code in modern programming languages (e.g. Python). This indicates that simple code modifications can result in significant performance gains when working with crowdsourced preferences.
We further evaluate Crowd-PrefPPO on a multi-objective robotics task, showing that Crowd-PrefPPO can identify minority groups in the crowd and demonstrating the method’s potential to learn minority reward functions (Figure 5). How to best incorporate this knowledge into the agent training loop remains an open question; our framework could lend itself to multiple approaches, e.g. learning a utility function that attempts to balance the multiple users’ desired behavior, learning personalized reward functions for specific user groups, or allowing human users to assign weights to each reward function. In future work, we hope to explore how to robustly account for multiple learned objectives across these various scenarios.
Several limitations of our approach are important to discuss. First, each user in the crowd must provide a preference label for each preference query. Future work will focus on the possibility of using Crowd-PrefRL when user feedback is sparse. In addition, this study uses only simulated human feedback, which allows for a systematic evaluation of Crowd-PrefRL and an analysis of its properties in a controlled setting. Future work will evaluate this method in a human subjects study.
Supplementary Materials
The following content was not necessarily subject to peer review.
Full Crowd-PrefRL Algorithm
The full Crowd-PrefRL algorithm is specified in Algorithm 1, where red text indicates differences from the framework in  Lee et al. (2021a).
@@ 6 Discussion
22:         $\mathcal{D}\leftarrow\mathcal{D}\cup\{(A,B,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\bm{y}}})\}$
@@
24:              Sample batch $\{(A,B,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\bm{y}}})_{j}\}_{j=1}^{|\mathcal{D}|}\sim\mathcal{D}$
@@
Experiment Setup Hyperparameters
@@
Our experiments use the Walker-walk, Quadruped-walk, and Cheetah-run environments from the DMControl suite (Tassa et al., 2018), as two of these environments (Walker-walk and Quadruped-walk) were previously used in Lee et al. (2021a). All hyperparameters for our method can be found in Table 1, and notably, are the same as reported in Lee et al. (2021a). Hyperparameters that we use for the MO-Hopper environment are the same as those used for the Hopper environment in stable-baselines3.